A boy shoves his stuffed toy zebra, which has mass
Answer #1

Initial kinetic energy and potential energy of the toy is lost to friction while sliding across the rough horizontal surface

KE+ PE = W friction

KE; + PE; = F friction *d

Frictional force = coefficient of friction * Normal reaction

KE; + PE; = 4N*d

mV? + mgD = umg *d

Cancel mass m

p*6rt = 76 +247

Divide all terms with \mu g

vo 2μg 9D μg μgd μg

2μg μ

ANSWER: {\color{Red} d=\frac{V^{2}}{2\mu g}+\frac{D}{\mu }}

==========================